# KiDS-Legacy: angular galaxy clustering from deep surveys with complex selection effects

### v1

@@
Ziang Yan (颜子昂) E-mail:yanza21@astro.rub.de11    Angus H. Wright 11    Nora Elisa Chisari 22    Christos Georgiou 22    Shahab Joudaki 3344    Arthur Loureiro 5566    Robert Reischke 77    Marika Asgari 88    Maciej Bilicki 99    Andrej Dvornik 11    Catherine Heymans 101011    Hendrik Hildebrandt 11    Priyanka Jalan 99    Benjamin Joachimi $1$    Giorgio Francesco Lesci 12121313    Shun-Sheng Li (李顺生) 14141515    Laila Linke 1616    Constance Mahony 1717181811    Lauro Moscardini 121213131919    Nicola R. Napolitano 2020    Benjamin Stölzner 11    Maximilian Von Wietersheim-Kramsta 21212222    Mijin Yoon 1414
Photometric galaxy surveys, despite their limited resolution along the line of sight, encode rich information about the large-scale structure (LSS) of the Universe thanks to the large number density and extensive depth of the data. However, the complicated selection effects in wide and deep surveys will potentially cause significant bias in the angular two-point correlation function (2PCF) measured from those surveys. In this paper, we measure the 2PCF from the newly published KiDS-Legacy sample. Given an $r$-band $5\sigma$ magnitude limit of $24.8$ and survey footprint of $1347$ deg2, it achieves an excellent combination of sky coverage and depth for such a measurement. We find that complex selection effects, primarily induced by varying seeing, introduce over-estimation of the 2PCF by approximately an order of magnitude. To correct for such effects, we apply a machine learning-based method to recover an “organised random” (OR) that presents the same selection pattern as the galaxy sample. The basic idea is to find the selection-induced clustering of galaxies using a combination of self-organising maps (SOM) and hierarchical clustering (HC). This unsupervised machine learning method is able to recover complicated selection effects without specifying their functional forms. We validate this “SOM+HC” method on mock deep galaxy samples with realistic systematics and selections derived from the KiDS-Legacy catalogue. Using mock data, we demonstrate that the OR delivers unbiased 2PCF cosmological parameter constraints, removing the $27\sigma$ offset in the galaxy bias parameter that is recovered when adopting uniform randoms. Blinded measurements on the real KiDS-Legacy data show that the corrected 2PCF is robust to the SOM+HC configuration near the optimal setup suggested by the mock tests. Our software is open-source for future usage.
@@ 1 Introduction
The fluctuation patterns of the mass distribution in the Universe, i.e., the large-scale structure (LSS), have been one of the central topics of modern cosmology (see, for example, Dodelson & Schmidt, 2020, for a comprehensive introduction to the LSS). Because the matter in the Universe is mainly made up of invisible dark matter, one needs to observe visible LSS tracers to infer the matter distribution. Among observations of various LSS tracers, large-scale galaxy surveys have provided robust constraints on cosmological parameters. Spectroscopic surveys such as the Sloan Digital Sky Survey (SDSS-V, Kollmeier et al., 2019) and the Dark Energy Spectroscopic Instrument (DESI, Levi et al., 2019) map 3-D galaxy distributions from which they measure baryon acoustic oscillations (BAOs) and redshift-space distortions (RSD) to constrain the Hubble constant, matter density, dark energy equation of state, and neutrino mass (Cuceu et al., 2019; Loureiro et al., 2019; Neveux et al., 2020). Recent photometric surveys, including the Kilo-Degree Survey (KiDS, Asgari et al., 2021), the Dark Energy Survey (DES, Amon et al., 2022; Secco et al., 2022), and the Hyper Suprime-Cam Subaru Strategic Program (HSC-SSP, Li et al., 2023b; Dalal et al., 2023) have obtained consistent joint constraints on the matter density, $\Omega_{\mathrm{m}}$ and fluctuation amplitude, $\sigma_{8}$, of the density field with a precision of $\sim 5\%$. As the precision improves, tensions between the probes have begun to emerge, an important one being the tension in $S_{8}$, the matter clustering amplitude parameter, between the CMB and the late-time Universe probes. (for a review, see Abdalla et al., 2022). Future surveys, including the Legacy Survey of Space and Time (LSST) at Vera Rubin Observatory (LSST Science Collaboration et al., 2009), the Euclid Survey (Euclid Collaboration et al., 2024), the Xuntian (Gong et al., 2019, also known as the Chinese Space Station Telescope, CSST) are promising to tackle these problems with better data quality and systematics estimations.
@@
The galaxy distribution encodes rich information about the LSS of the Universe and galaxy formation, so it has long been used as a tracer to study the LSS (Peebles, 1973; Shanks et al., 1983; Maddox et al., 1990; Efstathiou et al., 1990; Maddox et al., 1996; Nichol, 2007). On large scales, it can be estimated to be linearly biased relative to the matter distribution(see Desjacques et al., 2018, for a review on galaxy bias). More subtle bias terms can be modelled via an effective field theory (Baumann et al., 2012; Carrasco et al., 2012). On small scales, the galaxy distribution can be described by a halo occupation distribution model (HOD, Zheng et al., 2005) in the context of the halo model (Peacock & Smith, 2000; Seljak, 2000; Cooray & Sheth, 2002; Asgari et al., 2023). An important summary statistic of the galaxy distribution is the two-point correlation function (2PCF), which describes the galaxy clustering between two points as a function of their separations. The measurements of 2PCF have been used to constrain the matter density and clustering amplitude of the LSS since the 1980s (Davis & Peebles, 1983). Recently, a combination of galaxy clustering, galaxy-galaxy lensing, and cosmic shear, namely the $3\times 2$pt analysis, has greatly improved the constraining power on $\Omega_{\mathrm{m}}$ and $S_{8}$ (Heymans et al., 2021; Abbott et al., 2022; Sugiyama et al., 2023; Dvornik et al., 2023).
@@
Measurements of 3-dimensional galaxy clustering depend on reliable galaxy redshift measurements, typically carried out by spectroscopic observations that are generally time-consuming and limited in depth. While photometric surveys can reach deeper, they cannot obtain precise redshifts for each galaxy. In this case, given the angular coordinates of galaxy positions, one can measure the angular galaxy clustering with the 2PCF with respect to angular separation. With the calibrated redshift distribution of a photometric sample, the angular 2PCF can be modelled as projected 3-D 2PCF weighted by the redshift distribution. Angular 2PCF also encodes a significant amount of cosmological information (see, for example, Coupon et al., 2012) and is relatively easy to measure. However, several systematic uncertainties need to be considered to obtain accurate constraints. In terms of modelling, galaxy distributions are affected by redshift space distortions (RSDs, Kaiser, 1987) and cosmic magnification (Menard, 2002). On the observational side, galaxies are subject to selection effects by various observational systematics, including seeing, sky background, instrument response, survey strategy, galactic extinction, and so on. In practice, the selection effects are anisotropic given the various observational conditions and survey strategies, such that the galaxy sample will show a “variable depth” pattern that will introduce additional non-cosmological correlations that bias the angular 2PCF measurement and subsequent cosmological constraints [^1] This work primarily deals with angular 2PCF, so in the following text, if not explicitly stated otherwise, 2PCF means angular two-point correlation function..
@@
Figure 1: A flowchart illustrating the SOM+HC method to recover OR and correct selection effects in the 2PCF. Starting from the top: 1. input systematics; 2. SOM training: left panel: SOM grid colour-coded by galaxy number in each cell; right panel: projection of systematics vectors (grey dots) and SOM cells (red dots connecting) on the plane of two systematics. The projected adjacent SOM cells are connected with black lines; 3. HC output: left panel: SOM cells colour-coded according to hierarchical cluster indices; right panel: systematics vectors and weight vectors colour-coded by corresponding cluster indices; 4. effective areas corresponding to galaxies from each cluster; 5. recovered OR weight map which will be used in a subsequent 2PCF measurement. The details of this process are introduced in Section 4.
@@
  • Template-based corrections: this method parametrises the bias from selection effects in galaxy statistics and attempts to mitigate those effects with systematics templates (Weaverdyck & Huterer, 2021). Two methods are usually used: 1) template subtraction method (Ho et al., 2012; Ross et al., 2011) which models contamination terms as templates multiplied by an amplitude factor. After determining the amplitude factor, these terms are subtracted from the statistics. 2) mode projection (Leistedt et al., 2013; Leistedt & Peiris, 2014; Nicola et al., 2020; Berlfein et al., 2024): this method mitigates systematics contamination by marginalizing the functional contribution of systematic templates.
@@
  • Forward modelling with synthetic objects: this approach calibrates the observational selections by injecting artificial galaxies into realistic images (Bergé et al., 2013; Suchyta et al., 2016). The injected objects are passed through a realistic measurement process, resulting in a random catalogue mimicking realistic selection functions. This method is computationally demanding, but it is under active development for ongoing and future surveys (Everett et al., 2022; Kong et al., 2024).
@@
  • Machine-learning reconstruction of selection functions: this method aims at probing the relationship between observed galaxy number densities and multivariate systematics vectors, thus recovering the selection functions. For example, Rezaie et al. (2020) applies a deep neural network to recover the selection weights for galaxies from the Dark Energy Camera Legacy Survey (DECaLS, Dey et al., 2019), while Morrison & Hildebrandt (2015) applied the $k$-means algorithm on the CFHTLenS to find selection-induced galaxy clustering in high-dimensional systematics space.
Template-based correction methods depend on the functional assumptions of systematic contamination. For example, Nicola et al. (2020) assumes additive systematics contamination in the galaxy overdensity and tries to fit the linear factor from systematics maps. This method works when the contamination level is low (as in that work), but will potentially fail when the selection functions are complicated and correlated. On the other hand, forward modelling and machine-learning-based methods try to recover the overall selection effects without modelling the selection function of each systematic. Therefore, these methods are capable of capturing complicated selection effects, but they are valid only when the sample is large enough for the algorithm to learn the selection pattern.
@@
In practice, 2PCF can be estimated with the Landy-Szalay estimator (Landy & Szalay, 1993), which employs a random catalogue to factor out non-cosmological correlations. If the catalogue has significant selection effects, a uniform random catalogue (UR) will fail to capture them and prevent a bias in the estimated 2PCF. An “organised random” (OR), namely a tailored random catalogue reflecting the same selection effects, can be used instead to correct such biases. Johnston et al. (2021) proposed a machine-learning-based method to recover the OR. It uses a combination of self-organising maps and hierarchical clustering (SOM+HC) to identify galaxy clusters [^2] Note that here the galaxy clusters are in the systematics space, and are not related to gravitationally bounded galaxy clusters in the Universe. In the following text, we use the term cluster to refer to the cluster of galaxies in the systematics space. in the systematics space. Galaxies from the same cluster are assumed to be selected uniformly, so a UR with the same galaxy number is generated in the sky region occupied by those galaxies. The OR is constructed by combining the URs corresponding to all the clusters. Figure 1 provides a brief flowchart of this method which has been validated in Johnston et al. (2021) for the KiDS-1000 bright sample.
@@
In this work, we will apply the SOM+HC method to correct for complex selection effects in the full KiDS-Legacy sample. We will first present the SOM+HC methodology, then test it on mock galaxy samples generated from the Generator for Large-Scale Structure (GLASS, Tessore et al., 2023). The mock systematics and selections are applied based on the real KiDS-Legacy sample. We evaluate the method by comparing the 2PCF measured with the recovered OR with an unbiased 2PCF with no selection effect. Then we perform a preliminary, blinded 2PCF measurement from the real KiDS-Legacy catalogue. This work focuses on comparing and modelling the 2PCFs on linear scales, as most photometric surveys use such scales due to limitations in the theoretical modelling (Crocce et al., 2015; Rodríguez-Monroy et al., 2022). In forthcoming work, the OR recovered from the KiDS-Legacy catalogue will be used to measure the 2PCF for the KiDS-Legacy $6\times 2$-pt analysis, and will also be useful for future LSS surveys.
@@
The paper is structured as follows: Section 2 describes the definition and measurements of galaxy 2PCF; Section 3 presents the data and mock that we use to run and test the method; Section 4 introduces the SOM+HC method to correct the selection effects; Section 5 deals with validations of the SOM+HC methods; Section 6 presents a blinded measurement with the KiDS-legacy catalogue; and Section 7 summarises the conclusion of this work and discusses the advantages of our method and future prospects. Throughout this study, we assume a flat $\Lambda$CDM cosmology with the fixed cosmological parameters from Planck Collaboration (2020) as our background: $(h,\Omega_{\mathrm{c}}h^{2},\Omega_{\mathrm{b}}h^{2},\sigma_{8},n_{\mathrm{s}})=(0.676,0.119,0.022,0.81,0.967)$. The code used in this work is published as the tiaogeng333{CJK}UTF8gbsn Tiaogeng (调羹) is the Chinese word for “spoon”, more commonly used in southern China. It contains two characters: “tiao (调)” meaning to reconcile and “geng (羹)” referring to a Chinese-style thick soup. The code reconciles the unevenly observed sky, just as a tiaogeng stirs soup to make it taste more balanced and delicious. package [^4] https://github.com/yanzastro/tiaogeng/ for future usage.
@@ 2.1 Definition and connection with matter distribution
Galaxy 3-D 2PCFs describe the correlation of galaxy number overdensity between points separated by a certain distance. For two galaxy samples $a$ and $b$, their 2PCF is defined as (Peebles, 1973):
@@
$\delta_{\mathrm{g}}(\boldsymbol{x}):=\frac{n(\boldsymbol{x})-\bar{n}}{\bar{n}}\,,$ (2)
where $n(\boldsymbol{x})$ denotes the galaxy number density at $\boldsymbol{x}$ and $\bar{n}$ is the mean galaxy number density. Since we mainly focus on the galaxy 2PCF, we omit the subscript $g$ in the following formulas. We note to the reader that the 2PCF can also be defined for fluctuations of any LSS fields.
According to Parseval’s theorem, the 2PCF is the Fourier transform of the power spectrum, the correlation function in Fourier space. They are both widely used summary statistics of the spatial distribution of cosmological fields, and they both depend on the evolution and ingredients of the Universe. For detailed mathematical descriptions of the 2PCF and power spectra, we refer the reader to Dodelson & Schmidt (2020).
For photometric surveys, we usually cannot measure the distance of an individual galaxy. Thus we can only project galaxies onto the celestial sphere and study their angular distribution. We can define the angular two-point correlation function by replacing $\boldsymbol{x}$ with a 2-dimensional angular vector $\boldsymbol{\theta}$. Since we only care about the angular galaxy distribution in this paper, we use 2PCF to denote the angular two-point correlation function unless otherwise stated. Under the “flat-sky approximation”, the angular 2PCF is related to the angular power spectrum via (Peebles, 1973):
@@
where $J_{0}(x)$ is the zeroth order Bessel function of the first kind; $C^{ab}_{\ell}$ is the angular galaxy cross-correlation power spectrum of samples $a$ and $b$. Note that this formula works for 2PCF and angular power spectra between any cosmological fields. At sufficiently large $\ell$, the angular power spectra can be calculated via the Limber approximation (Limber, 1953):
$C^{ab}_{\ell}=\int_{0}^{\chi_{\mathrm{H}}}\frac{{\mathrm{d}}\chi}{\chi^{2}}W^{a}(\chi)W^{b}(\chi)P^{ab}_{\mathrm{gg}}\left(\frac{\ell+1/2}{\chi},z(\chi)\right)\,,$ (4)
where $\chi$ is the comoving distance and the subscript $\rm H$ is the horizon; $P_{\mathrm{gg}}(k,z)$ is the galaxy power spectra at redshift $z$; $W^{a}(\chi)$ and $W^{b}(\chi)$ are the radial kernels of the two galaxy samples. Note that this formulation works for flat cosmology. For sample $a$, given its normalised galaxy redshift distributions $n^{a}(z)$, its radial kernel is
@@
The galaxy power spectra $P_{\mathrm{gg}}(k,z)$ encode the 3-D fluctuations of galaxy distributions, which follow the fluctuations of underlying mass distribution. On small scales, the non-linear clustering of galaxies can be modelled as a halo model-based halo occupation distribution (HOD, Seljak, 2000; Cooray & Sheth, 2002; Zheng et al., 2005). On large scales, galaxy fluctuations can be approximated as linearly biased matter fluctuations with $\delta_{\mathrm{g}}(\boldsymbol{x},z)=b_{\mathrm{g}}(z)\delta_{\mathrm{m}}(\boldsymbol{x},z)$ where $b_{\mathrm{g}}(z)$ is the redshift-dependent galaxy bias. Thus $P^{ab}_{\mathrm{gg}}(k,z)=b^{a}_{\mathrm{g}}(z)b^{b}_{\mathrm{g}}(z)P_{\mathrm{m}}(k,z)$. We can take the galaxy biases outside of the Limber integral by approximating them at the mean redshift $\bar{z}^{a}\equiv\int n^{a}(z)z{\mathrm{d}}z/\int n^{a}(z){\mathrm{d}}z$ of the galaxy samples (denoted as $b^{a}\equiv b_{\mathrm{g}}(\bar{z}^{a})$ and $b^{b}\equiv b_{\mathrm{g}}(\bar{z}^{b})$). Therefore we have, on linear scales:

### v2

@@
Ziang Yan E-mail:yanza21@astro.rub.de11    Angus H. Wright 11    Nora Elisa Chisari 22 14 14    Christos Georgiou 22    Shahab Joudaki 3344    Arthur Loureiro 5566    Robert Reischke 77    Marika Asgari 88    Maciej Bilicki 99    Andrej Dvornik 11    Catherine Heymans 101011    Hendrik Hildebrandt 11    Priyanka Jalan 99    Benjamin Joachimi $1$    Giorgio Francesco Lesci 12121313    Shun-Sheng Li 14141515    Laila Linke 1616    Constance Mahony 1717181811    Lauro Moscardini 121213131919    Nicola R. Napolitano 2020    Benjamin Stölzner 11    Maximilian Von Wietersheim-Kramsta 21212222    Mijin Yoon 1414
Photometric galaxy surveys, despite their limited resolution along the line of sight, encode rich information about the large-scale structure (LSS) of the Universe thanks to the high number density and extensive depth of the data. However, the complicated selection effects in wide and deep surveys can potentially cause significant bias in the angular two-point correlation function (2PCF) measured from those surveys. In this paper, we measure the 2PCF from the newly published KiDS-Legacy sample. Given an $r$-band $5\sigma$ magnitude limit of $24.8$ and survey footprint of $1347$ deg2, it achieves an excellent combination of sky coverage and depth for such a measurement. We find that complex selection effects, primarily induced by varying seeing, introduce over-estimation of the 2PCF by approximately an order of magnitude. To correct for such effects, we apply a machine learning-based method to recover an organised random (OR) that presents the same selection pattern as the galaxy sample. The basic idea is to find the selection-induced clustering of galaxies using a combination of self-organising maps (SOMs) and hierarchical clustering (HC). This unsupervised machine learning method is able to recover complicated selection effects without specifying their functional forms. We validate this SOM+HC method on mock deep galaxy samples with realistic systematics and selections derived from the KiDS-Legacy catalogue. Using mock data, we demonstrate that the OR delivers unbiased 2PCF cosmological parameter constraints, removing the $27\sigma$ offset in the galaxy bias parameter that is recovered when adopting uniform randoms. Blinded measurements on the real KiDS-Legacy data show that the corrected 2PCF is robust to the SOM+HC configuration near the optimal set-up suggested by the mock tests.
@@ 1 Introduction
The fluctuation patterns of the mass distribution in the Universe (i.e. the large-scale structure, LSS) is one of the central topics of modern cosmology (see, for example, Dodelson & Schmidt, 2020, for a comprehensive introduction to the LSS). Because the matter in the Universe is mainly made up of invisible dark matter, one needs to observe visible LSS tracers to infer the matter distribution. Among observations of various LSS tracers, large-scale galaxy surveys have provided robust constraints on cosmological parameters. Spectroscopic surveys such as the Sloan Digital Sky Survey (SDSS-IV; Alam et al., 2021) and the Dark Energy Spectroscopic Instrument (DESI; DESI Collaboration et al., 2024) map 3D galaxy distributions from which they measure baryon acoustic oscillations (BAOs) and redshift-space distortions (RSDs) to constrain the Hubble constant, matter density, dark energy equation of state, and neutrino mass (Cuceu et al., 2019; Loureiro et al., 2019; Neveux et al., 2020). Recent photometric surveys, including the Kilo-Degree Survey (KiDS; Asgari et al., 2021), the Dark Energy Survey (DES; Amon et al., 2022; Secco et al., 2022), and the Hyper Suprime-Cam Subaru Strategic Program (HSC-SSP; Li et al., 2023b; Dalal et al., 2023) have obtained consistent joint constraints on the matter density, $\Omega_{\mathrm{m}}$, and fluctuation amplitude, $\sigma_{8}$, of the density field with a precision of $\sim 5\%$. As the precision improves, tensions between the probes have begun to emerge, an important one being the tension in $S_{8}$, the matter clustering amplitude parameter, between the CMB and the late-time Universe probes. (for a review, see Abdalla et al., 2022). Future surveys, including the Legacy Survey of Space and Time (LSST) at the Vera Rubin Observatory (LSST Science Collaboration et al., 2009), the Euclid Survey (Euclid Collaboration et al., 2024), the Xuntian (Gong et al., 2019, also known as the Chinese Space Station Telescope, CSST) are promising to tackle these problems with better data quality and systematics estimations.
@@
The galaxy distribution encodes rich information about the LSS of the Universe and galaxy formation, so it has long been used as a tracer to study the LSS (Peebles, 1973; Shanks et al., 1983; Maddox et al., 1990; Efstathiou et al., 1990; Maddox et al., 1996; Nichol, 2007). On large scales, it can be estimated to be linearly biased relative to the matter distribution(see Desjacques et al., 2018, for a review on galaxy bias). Subtler bias terms can be modelled via an effective field theory (Baumann et al., 2012; Carrasco et al., 2012). On small scales, the galaxy distribution can be described by a halo occupation distribution model (HOD; Zheng et al., 2005) in the context of the halo model (Peacock & Smith, 2000; Seljak, 2000; Cooray & Sheth, 2002; Asgari et al., 2023). An important summary statistic of the galaxy distribution is the two-point correlation function (2PCF), which describes the galaxy clustering between two points as a function of their separations. The measurements of 2PCF have been used to constrain the matter density and clustering amplitude of the LSS since the 1980s (Davis & Peebles, 1983). Recently, a combination of galaxy clustering, galaxy-galaxy lensing, and cosmic shear, called the $3\times 2$pt analysis, has greatly improved the constraining power on $\Omega_{\mathrm{m}}$ and $S_{8}$ (Heymans et al., 2021; Abbott et al., 2022; Sugiyama et al., 2023; Dvornik et al., 2023).
@@
Measurements of three-dimensional galaxy clustering depend on reliable galaxy redshift measurements, typically carried out by spectroscopic observations that are generally time-consuming and limited in depth. While photometric surveys can reach deeper, they cannot obtain precise redshifts for each galaxy. In this case, given the angular coordinates of galaxy positions, one can measure the angular galaxy clustering with the 2PCF with respect to angular separation. With the calibrated redshift distribution of a photometric sample, the angular 2PCF can be modelled as projected 3-D 2PCF weighted by the redshift distribution. Angular 2PCF also encodes a significant amount of cosmological information (see, for example; Coupon et al., 2012) and is relatively easy to measure. However, several systematic uncertainties need to be considered to obtain accurate constraints. In terms of modelling, galaxy distributions are affected by redshift space distortions (RSDs; Kaiser, 1987) and cosmic magnification (Menard, 2002). On the observational side, galaxies are subject to selection effects by various observational systematics, including seeing, sky background, instrument response, survey strategy, galactic extinction, and so on. In practice, the selection effects are anisotropic given the various observational conditions and survey strategies, such that the galaxy sample will show a variable depth pattern that will introduce additional non-cosmological correlations that bias the angular 2PCF measurement and subsequent cosmological constraints [^1] This work primarily deals with angular 2PCF, so in the following text, if not explicitly stated otherwise, 2PCF means angular two-point correlation function..
@@
  • Mitigating selection bias in statistics. Correcting the selection effects with the following two methods (Weaverdyck & Huterer, 2021): 1) template subtraction method (Ho et al., 2012; Ross et al., 2011), which models contamination terms as templates multiplied by an amplitude factor. After determining the amplitude factor, these terms are subtracted from the statistics; 2) mode projection (Leistedt et al., 2013; Leistedt & Peiris, 2014; Nicola et al., 2020; Berlfein et al., 2024), which mitigates systematics contamination by marginalising the functional contribution of systematic templates.
@@
  • Forward modelling with synthetic objects. This approach calibrates the observational selections by injecting artificial galaxies into realistic images (Bergé et al., 2013; Suchyta et al., 2016). The injected objects are passed through a realistic measurement process, resulting in a random catalogue mimicking realistic selection functions. This method is computationally demanding, but it is under active development for ongoing and future surveys (Everett et al., 2022; Kong et al., 2024).
  • Reconstructing the selection function by regression. This method aims to probe the relationship between observed galaxy number densities and systematics values to construct weight maps or random catalogues that reflect the reconstructed selection functions. For example, Elvin-Poole et al. (2018), Rezaie et al. (2020), and Rodríguez-Monroy et al. (2022) regress galaxy number and systematics with the help of systematics templates, while Morrison & Hildebrandt (2015) accounts for selection effects by finding selection-induced galaxy clustering in high-dimensional systematics space.
@@
A subclass of regression methods uses machine learning to learn the relationship between galaxy number densities and systematics values. This class of methods is more capable of finding complex, non-linear, and potentially correlated selection effects. For example, Rezaie et al. (2020) uses deep neural networks to derive weights of galaxies selected from the Dark Energy Camera Legacy Survey (DECaLS; Dey et al., 2019); Morrison & Hildebrandt (2015) applied a $k$-means clustering method in the high-dimensional density-systematics space to create weight maps. The method used in this work belongs to this category.
@@
We focus on 2PCF in this work. In practice, 2PCF can be estimated with the Landy-Szalay estimator (Landy & Szalay, 1993), which employs a random catalogue to factor out non-cosmological correlations. If the catalogue has significant selection effects, a uniform random catalogue (UR) will fail to capture them and will not prevent a bias in the estimated 2PCF. An organised random catalogue (OR) is a tailored random catalogue reflecting the same selection effects that can be used to correct such biases. Johnston et al. (2021) proposed a machine-learning-based method to recover the OR. It uses a combination of self-organising maps and hierarchical clustering (SOM+HC) to identify galaxy clusters [^2] Here the galaxy clusters are in the systematics space, and are not related to gravitationally bounded galaxy clusters in the Universe. In the following text we use the term cluster to refer to the cluster of galaxies in the systematics space. in the systematics space. Galaxies from the same cluster are assumed to be selected uniformly, so a UR with the same galaxy number is generated in the sky region occupied by those galaxies. The OR is constructed by combining the URs corresponding to all the clusters. This method has been validated in Johnston et al. (2021) for the KiDS-1000 bright sample.
@@
In this work, we apply the SOM+HC method to correct for complex selection effects in the full KiDS-Legacy sample. We first present the SOM+HC methodology, then test it on mock galaxy samples generated from the Generator for Large-Scale Structure (GLASS; Tessore et al., 2023). The mock systematics and selections are applied based on the real KiDS-Legacy sample. We evaluate the method by comparing the 2PCF measured with the recovered OR with an unbiased 2PCF with no selection effect. Then we perform a preliminary, blinded 2PCF measurement from the real KiDS-Legacy catalogue. This work focuses on comparing and modelling the 2PCFs on linear scales, as most photometric surveys use such scales due to limitations in the theoretical modelling (Crocce et al., 2015; Rodríguez-Monroy et al., 2022). In forthcoming work, the OR recovered from the KiDS-Legacy catalogue will be used to measure the 2PCF for the KiDS-Legacy $6\times 2$-pt analysis, and will also be useful for future LSS surveys.
@@
The paper is structured as follows: Sect. 2 describes the definition and measurements of galaxy 2PCF; Sect. 3 presents the data and mock that we use to run and test the method; Sect. 4 introduces the SOM+HC method to correct the selection effects; Sect. 5 deals with validations of the SOM+HC methods; Sect. 6 presents a blinded measurement with the KiDS-legacy catalogue; and Sect. 7 summarises the conclusion of this work and discusses the advantages of our method and future prospects. Throughout this study we assume a flat $\Lambda$CDM cosmology with the fixed cosmological parameters from Planck Collaboration (2020) as our background: $(h,\Omega_{\mathrm{c}}h^{2},\Omega_{\mathrm{b}}h^{2},\sigma_{8},n_{\mathrm{s}})=(0.676,0.119,0.022,0.81,0.967)$. The code used in this work is published as the tiaogeng333{CJK}UTF8gbsn Tiaogeng (调羹) is the Chinese word for ‘spoon’, more commonly used in southern China. It contains two characters: tiao (调) meaning to reconcile and geng (羹) referring to a Chinese-style thick soup. The code reconciles the unevenly observed sky, just as a tiaogeng stirs soup to make it taste more balanced and delicious. package [^4] https://github.com/yanzastro/tiaogeng/ for future use.
@@ 2.1 Definition and connection with matter distribution
Galaxy 3-D 2PCFs describe the correlation of galaxy number over-density between points separated by a certain distance. For two galaxy samples $a$ and $b$, their 2PCF is defined as (Peebles, 1973)
@@
$\delta_{\mathrm{g}}(\boldsymbol{x}):=\frac{n(\boldsymbol{x})-\bar{n}}{\bar{n}}\,.$ (2)
Here $n(\boldsymbol{x})$ denotes the galaxy number density at $\boldsymbol{x}$ and $\bar{n}$ is the mean galaxy number density. Since we mainly focus on the galaxy 2PCF, we omit the subscript $g$ in the following formulas. We note that the 2PCF can also be defined for fluctuations of any LSS fields.
According to Parseval’s theorem, the 2PCF is the Fourier transform of the power spectrum, the correlation function in Fourier space. They are both widely used summary statistics of the spatial distribution of cosmological fields, and they both depend on the evolution and ingredients of the Universe. For detailed mathematical descriptions of the 2PCF and power spectra, we refer to Dodelson & Schmidt (2020).
For photometric surveys, we usually cannot measure the distance of an individual galaxy. Thus we can only project galaxies onto the celestial sphere and study their angular distribution. We can define the angular two-point correlation function by replacing $\boldsymbol{x}$ with a 2-dimensional angular vector $\boldsymbol{\theta}$. Since we only care about the angular galaxy distribution in this paper, we use 2PCF to denote the angular two-point correlation function unless otherwise stated. Under the ‘flat-sky approximation’, the angular 2PCF is related to the angular power spectrum via (Peebles, 1973)
@@
where $J_{0}(x)$ is the zeroth order Bessel function of the first kind; $C^{ab}_{\ell}$ is the angular galaxy cross-correlation power spectrum of samples $a$ and $b$. We note that this formula works for 2PCF and angular power spectra between any cosmological fields. At sufficiently large $\ell$, the angular power spectra can be calculated via the Limber approximation (Limber, 1953):
$C^{ab}_{\ell}=\int_{0}^{\chi_{\mathrm{H}}}\frac{{\mathrm{d}}\chi}{\chi^{2}}W^{a}(\chi)W^{b}(\chi)P^{ab}_{\mathrm{gg}}\left(\frac{\ell+1/2}{\chi},z(\chi)\right)\,.$ (4)
Here $\chi$ is the comoving distance and the subscript $\rm H$ is the horizon; $P_{\mathrm{gg}}(k,z)$ is the galaxy power spectra at redshift $z$; $W^{a}(\chi)$ and $W^{b}(\chi)$ are the radial kernels of the two galaxy samples. We note that this formulation works for flat cosmology. For sample $a$, given its normalised galaxy redshift distributions $n^{a}(z)$, its radial kernel is
@@
The galaxy power spectra $P_{\mathrm{gg}}(k,z)$ encode the 3-D fluctuations of galaxy distributions, which follow the fluctuations of underlying mass distribution. On small scales, the non-linear clustering of galaxies can be modelled as a halo model-based halo occupation distribution (HOD; Seljak, 2000; Cooray & Sheth, 2002; Zheng et al., 2005). On large scales, galaxy fluctuations can be approximated as linearly biased matter fluctuations with $\delta_{\mathrm{g}}(\boldsymbol{x},z)=b_{\mathrm{g}}(z)\delta_{\mathrm{m}}(\boldsymbol{x},z)$ where $b_{\mathrm{g}}(z)$ is the redshift-dependent galaxy bias. Thus $P^{ab}_{\mathrm{gg}}(k,z)=b^{a}_{\mathrm{g}}(z)b^{b}_{\mathrm{g}}(z)P_{\mathrm{m}}(k,z)$. We can take the galaxy biases outside of the Limber integral by approximating them at the mean redshift $\bar{z}^{a}\equiv\int n^{a}(z)z{\mathrm{d}}z/\int n^{a}(z){\mathrm{d}}z$ of the galaxy samples (denoted as $b^{a}\equiv b_{\mathrm{g}}(\bar{z}^{a})$ and $b^{b}\equiv b_{\mathrm{g}}(\bar{z}^{b})$). Therefore, on linear scales we have